OpenStudy (anonymous):

the algebraic equation for two savings plans are y=20x+400 and y=10x+600 solve the system of equations algebraically to determine when the two savings plans will yield the same balance and show all work

OpenStudy (anonymous):

Equate the 2 equations together. y=20x+400 y=10x+600 20x+400 = 10x+600 20x - 10x = 600- 400 10x = 200 x= 200/10 = 20.